OpenStudy (mendicant_bias):

I'm having trouble understanding an intermediate step in a solution of an indefinite integral, can somebody help me? (Integral below in a minute).

OpenStudy (mendicant_bias):

\[\int\limits_{?}^{?} \frac{3 }{ x + 2\sqrt{x} }\]

OpenStudy (mendicant_bias):

First, the constant of 3 is factored outside the integral, and then after that, by u-substitution, they get this:

OpenStudy (mendicant_bias):

\[6 \int\limits_{}^{} \frac{ u }{ u ^{2}+u }du\]

OpenStudy (mendicant_bias):

What I don't understand is why the constant (3) doubled, and how, when using u-substitution where u is (sqrt)x and du is 1/2(sqrt)x du, why the numerator of the substituted integral is u.

OpenStudy (mendicant_bias):

Let me clarify my thought process by showing what I would do, and then maybe somebody can point out where I'm wrong, give me a moment:

OpenStudy (mendicant_bias):

\[\int\limits_{}^{}\frac{ 3 }{ x+2\sqrt{x} } = 3\int\limits_{}^{}\frac{ 1 }{ x+2\sqrt{x} }\] \[u = \sqrt{x}, du = \frac{ 1 }{ 2\sqrt{x} }dx\] \[3\int\limits_{}^{}\frac{ 1 }{ u ^{2}+2u }du\] I would have ended up like that. I don't know what's wrong with my substitution. I don't understand how they're getting a u in their numerator, or why the constant factored out changes from three to six.

OpenStudy (anonymous):

To me it seems that you have a recipe to 'simplify' the integral before trying to solve it. Forget about the constant 3 (or 6) for a moment and focus on the replacement of x. As there is a root(x) in your formula, you would like to find an easy replacement which removes this root(x), which looks too scary to work with. The best choice is then: replace x by u^2, because root(x) will then be replaced by root(u^2) which happens to be just 'u'. In your original problem, there also needs to be a dx at the end of your integral. Since you have replaced x by u^2, you also need to replace dx by d(u^2), which is 2*du. This is where the 2 is introduced on the top part of your integral. Together with the 3, which was already there, you end up with 6 on the topside. This '6' can be brought before the integral sign as you already captured in your posting. Hope this is clear now ?

OpenStudy (anonymous):

Yep, in your approach you didn't include the dx in the original problem.

OpenStudy (psymon):

I do it and get everything except u^2 +u. I get the 6 and the u on top, but I get u^2 + 2u on bottom, lol. I have an extra 2 from what he said the answer was.

OpenStudy (mendicant_bias):

Sorry, I forget the dx's. They're obviously important but I have a habit of it seeming like a pedantic thing tacked on the bottom. One second, I'm having a little trouble understanding your explanation, NL, but thanks, and give me a minute.

OpenStudy (anonymous):

I think NL pointed it out (and not sure if you do it), but it may help to solve for dx when doing u-sub. I always take an extra moment to solve for it and sub it back into the problem and then simplify.

OpenStudy (psymon):

Yea, what you typed was an error. Where you put u^2 + u in the denominator, it's supposed to be u^2 + 2u.

OpenStudy (mendicant_bias):

Oh, got it! I got the answer, but I came to it in a slightly different way that still works repeatedly.